A body starting from rest moves with constant acceleration.the ratio of distance covered by the body during the 5th sec to that covered in 5 sec
 

Grade:11

2 Answers

chetan jangir
4 years ago
IT is very simple to deal with such problems. We start with basics we can write distance travelled in 5th second is equal to total covered distance in 5 seconds minus total covered distance in 4 seconds.we get 1/2a5^2 -1/2 a4^2( we have let acceleration  a  it will also be cancelled out in taking ratio) We get 4.5a .And as we know distance covered in first 5 seconds is 1/2a5^2 that is 12.5a. Taking ratio we get 9/25.Which is required  answer .
Note .We can generalise a formula for distance covered in nth second it comes u/2 +a/2(2n-1)  where u=initial  velocity a = constant acceleration. 
Saurabh Koranglekar
4 years ago
Dear student

S5 = 0.5a(25)

S5th = 0.5a( 25-16) = 0.5a(9)

Ratio s5th/ s5 = 9 /25

Regards